OpenStudy (anonymous):

State the quadrant in which θ lies. a) sin θ < 0 and tan θ >0 b) cos θ < 0 and sin θ < 0 c) tan θ > 0 and cos θ >0 d) tan θ < 0 and cos θ < 0 e) csc θ > 0 and cot θ < 0

OpenStudy (isaiah.feynman):

a) Third b) Third c)Third d)Second e)Second

OpenStudy (anonymous):

Hm. in C.. If \[ tan(\theta) > 0 \quad\quad \text{ and } \quad\quad cos(\theta) > 0 \] Since \[ \tan(\theta) = \frac{sin(\theta)}{cos(\theta)}\\ \text{Then } sin(\theta) \text{ ought to be positive} \] So if both the cos and the sin positive it should be first one not third. Unless I'm missing something

OpenStudy (isaiah.feynman):

@pitamar You are correct, its in the first one. Thanks for the heads up.